If a 50 lb force is applied at the end of a wrench that is 1 ft long, how much torque
is developed?
1
Expert's answer
2021-10-19T15:17:21-0400

Given:

"F=50\:\rm lb"

"d=1\:\rm ft"


The torque

"\tau=Fd\\=\rm (50\: lb)*(1\:ft)=50\: lb\cdot ft=68\: N\cdot m"